OpenStudy (anonymous):

Determine the largest acute angle of a right triangle which has legs of length 5 and 6 Determine the largest acute angle of a right triangle which has legs of length 5 and 6 @Mathematics

OpenStudy (anonymous):

Can yiou answer my ques.

OpenStudy (anonymous):

\[\frac{\text{ArcTan}\left[\frac{6}{5}\right]}{\pi }180=50.2{}^{\circ} \]